OpenStudy (anonymous):

Determinant with variable? A = [1-u -1 2] [0 4-u 3] [0 4 -u] Is it possible to get something like -u^3 +5u^2 -16u +12? If it's wrong, any help would be appreciated. Thanks!

OpenStudy (anonymous):

(1-u) * [(4-u+u) - (12) ] = 8u - 8

OpenStudy (anonymous):

isn't it (4-u)(u)?

OpenStudy (tkhunny):

Close. (4-u)(-u)

OpenStudy (anonymous):

opps my bad. Yea you're right

OpenStudy (anonymous):

so i did that and ended up with that polynomial of degree 3

OpenStudy (anonymous):

(1-u) * [(4-u)(-u) - (12) ] = -u^3+5 u^2+8 u-12. So yes, you're correct

OpenStudy (anonymous):

Thanks! also any advice for finding a value of u to make A invertible?

OpenStudy (anonymous):

-u^3+5 u^2+8 u-12 ≠ 0 (not equal) u ≠ -2, 1, 6